# arXiv:2512.09295v2[math.ST] 01 Mar 2026

### v2

@@ 1.3 Relations to the Literature
The closest works to ours are [GS24, JIS25, Lia25], which study denoisers motivated by OT theory [AGS05, Vil08]. Motivated by the fact that Bayes-optimal and empirical Bayes denoisers typically shrink the distribution too aggressively, these three papers propose different remedies. [GS24] and [JIS25] assume the existence of certain structural information of the prior distribution $P$, and solve for denoisers that minimize the mean squared error $\mathop{\mathbb{E}}\|T(Y)-X\|^{2}$ subject to such distributional constraints, say $T\sharp Q=P$; one can also view them as post-processing empirical Bayes denoisers to avoid over-shrinkage for better distributional matching. In contrast, this paper and [Lia25] do not require any a priori knowledge of $P$, and construct optimal transport-inspired denoisers solely based on the noisy measurements $Y\sim Q$. This paper generalizes the results in [Lia25] to arbitrary higher-order denoisers $T_{K},K\geq 1$ in the univariate setting. This paper characterizes the fundamental role of higher-order score functions in constructing the hierarchy of OT denoisers $T_{1},T_{2},\ldots,T_{\infty}$, and reveals the combinatorial structure underlying the infinite expansion of the optimal transport map via Bell polynomial recursions.
1.4 Notations
@@
We use the following notations throughout the paper. We use lower-case letters such as $x,y,z$ to denote scalars in $\mathbb{R}$, and upper-case letters such as $X,Y,Z$ denote random variables in $\mathbb{R}$. We reserve $f,g,h:\mathbb{R}\to\mathbb{R}$ to denote functions, and $F,G,T:\mathbb{R}\to\mathbb{R}$ to denote measurable maps. $P,Q$ denote probability distributions on $\mathbb{R}$, with density functions $p,q:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ respectively w.r.t. the Lebesgue measure; $\phi:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ denotes the standard normal density function. $\sigma>0$ denotes the noise level, and we call $\eta=\sigma^{2}/2$ the noise parameter.
@@
For a smooth function $f:\mathbb{R}\to\mathbb{R}$ and an integer $k$, we use $f^{(k)}$ to denote the $k$-th order derivative of $f$. For a probability distribution $P$ on $\mathbb{R}$ with density $p$, we use $\mathcal{L}(X)$ to denote the distribution of the random variable $X\sim P$. We use $T\sharp Q$ to denote the push-forward distribution of $Q$ via a measurable map $T:\mathbb{R}\to\mathbb{R}$, i.e., for any measurable set $A\subseteq\mathbb{R}$, $(T\sharp Q)(A)=P_{Y\sim Q}(T(Y)\in A)$. For two probability distributions $P,Q$ on $\mathbb{R}$, we use $W_{r}(P,Q)$ to denote the $r$-Wasserstein distance between $P$ and $Q$ with $r\geq 1$. Given a function $f$ and a probability distribution $Q$, we denote the $L^{2}(Q)$-norm as $\|f\|_{L^{2}(Q)}=\left([\mathop{\mathbb{E}}_{Y\sim Q}f(Y)^{2}]\right)^{1/2}$. We use $\|f\|_{\infty}=\sup_{x}|f(x)|$ to denote the $L^{\infty}$-norm of $f$, $\|f\|_{\mathrm{Lip}}$ to denote the uniform Lipschitz constant for $f$.
@@
Given i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$, we denote the empirical expectation $\widehat{\mathop{\mathbb{E}}}_{n}f=\tfrac{1}{n}\sum_{i=1}^{n}f(Y_{i})$. $\mathcal{F},\mathcal{H}$ denote function classes. We use $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|)$ to denote the covering number of a function class $\mathcal{F}$ at scale $\epsilon$ under the norm $\|\cdot\|$. We use the standard asymptotic notations $f(n)=O(g(n))$, or equivalently $f(n)\precsim g(n)$, if there exists a constant $C>0$ such that $|f(n)|\leq C|g(n)|$ for all sufficiently large $n$; we wrote $f(n)\asymp g(n)$ if $f(n)\precsim g(n)$ and $g(n)\precsim f(n)$.
@@ 2 Optimal Transport Expansion and Noise Asymptotics
We note that it is without loss of generality to assume $\sigma\in[0,1]$ and thus $\eta=\sigma^{2}/2\in[0,1/2]$: denote $\gamma^{2}:=\mathrm{Var}[Y]$, then we can rescale $Y$ to $\widetilde{Y}=Y/\gamma=\widetilde{X}+\widetilde{\sigma}Z$ where $\widetilde{X}=X/\gamma$ and $\widetilde{\sigma}=\sigma/\gamma\in[0,1]$. This illustrates that the noise level $\sigma$ can always be normalized to be at most $1$ by rescaling the data.
@@
Definition 1 (Model).
Let $X\in\mathbb{R}$ be a random variable with distribution $P$ and a valid density $p$ w.r.t. the Lebesgue measure. Let $Z\sim\mathcal{N}(0,1)$ be independent standard normal noise. Consider the noisy measurement $Y=X+\sigma Z$ with distribution $Q$ and density $q$. Let $\eta=\sigma^{2}/2$.
@@
[…]
The notation $F^{(k)},G^{(k)}$ denote the $k$-th order derivatives of $F,G$ respectively. $F^{-1},G^{-1}:[0,1]\to\mathbb{R}$ denote the quantile functions (generalized inverse) of $F$ and $G$ respectively, defined as
@@ 2.1 F-expansion of the Optimal Transport Map
Definition 2 (Optimal Transport Map).
@@
Such a map exists and is unique almost everywhere when $P,Q$ are absolutely continuous w.r.t the Lebesgue measure on $\mathbb{R}$ [Vil03, Theorem 2.18], and minimizes the $r$-Wasserstein distance between $P$ and $Q$ for any $r\geq 1$,
[…]
@@
Definition 3 (Bell Polynomials).
@@
$B_{n,k}(x_{1},x_{2},\ldots,x_{n-k+1})=\sum\frac{n!}{j_{1}!j_{2}!\cdots j_{n-k+1}!}\left(\frac{x_{1}}{1!}\right)^{j_{1}}\left(\frac{x_{2}}{2!}\right)^{j_{2}}\cdots\left(\frac{x_{n-k+1}}{(n-k+1)!}\right)^{j_{n-k+1}}\;,$
@@
Theorem 1 (F-expansion of the Optimal Transport Map).
Consider the additive Gaussian noise model in Definition 1. Assume that $F$ is infinitely differentiable. Define the following series expansion
@@
where $g_{1},g_{2},\ldots:\mathbb{R}\rightarrow\mathbb{R}$ are functions defined iteratively through the Bell polynomials $B_{n,k}$ in Definition 3:
@@
Then, for $y\in\mathbb{R}$ such that $T_{F}(y)$ is absolutely convergent, $T_{F}$ represents the optimal transport map pushing $Q$ onto $P$ as in Definition 2, namely,
@@
    Equation (1) reads $g_{2}=-\frac{F^{(2)}}{F^{(1)}}B_{2,2}(g_{1})+\frac{F^{(4)}}{F^{(1)}}$, and thus we have
@@
    Equation (1) reads $g_{3}=-\frac{F^{(2)}}{F^{(1)}}B_{3,2}(g_{1},g_{2})-\frac{F^{(3)}}{F^{(1)}}B_{3,3}(g_{1})+\frac{F^{(6)}}{F^{(1)}}$, and thus we have
@@
    $B_{4,1}(g_{1},g_{2},g_{3},g_{4})=g_{4},\quad B_{4,2}(g_{1},g_{2},g_{3})=4g_{1}g_{3}+3g_{2}^{2},\quad B_{4,3}(g_{1},g_{2})=6g_{1}^{2}g_{2},\quad B_{4,4}(g_{1})=g_{1}^{4}.$
    Equation (1) reads $g_{4}=-\frac{F^{(2)}}{F^{(1)}}B_{4,2}(g_{1},g_{2},g_{3})-\frac{F^{(3)}}{F^{(1)}}B_{4,3}(g_{1},g_{2})-\frac{F^{(4)}}{F^{(1)}}B_{4,4}(g_{1})+\frac{F^{(8)}}{F^{(1)}}$, and thus we have
@@ 2.2 Higher-Order Noise Asymptotics
In this section, we derive the accuracy of higher-order denoisers based on the noise asymptotic expansion of the optimal transport map in Theorem 1. Define the $K$-th order denoiser $T_{K}:\mathrm{supp}(P)\to\mathbb{R}$ as the truncated series expansion of $T$ up to the $K$-th order term,
@@
Given that $\{g_{i}\}_{i=1}^{K}$ depends on higher-order score functions $\tfrac{F^{(k)}}{F^{(1)}},k\leq 2K$ well-defined on the support of $P$, thus we only concern the behavior of $T_{K}$ on the support of $P$.
@@
[…]
In the case when $\mathrm{supp}(P)=\mathbb{R}$, this definition coincides with the standard $r$-Wasserstein distance $W_{r}(T\sharp Q,P)$; see Lemma 4 for this equivalence.
The next theorem studies the $W_{r}^{\ast}(\cdot,\cdot)$ between the denoised distribution $T_{K}\sharp Q$ and the signal distribution $P$, as well as the uniform approximation error between the denoiser $T_{K}$ and the optimal transport map $F^{-1}\circ G$ restricted on the support of $P$.
Theorem 2 (Higher-Order Accuracy).
Let $K\geq 1$ and consider the additive Gaussian noise model in Definition 1 with $F$ that is $(2K+2)$-differentiable. Let $T_{K}$ be the $K$-th order denoiser defined above with $\{g_{i}\}_{i=1}^{K}$ defined in Theorem 1. Assume that $\mathrm{supp}(P)$ is compact and $\inf_{x\in\mathrm{supp}(P)}F^{(1)}(x)>0$, and that $T_{K}$ is monotonically increasing. Recall (2.4), for any $r\geq 1$, we have
@@
Remark 2.
As noted earlier, we can assume that $\sigma\in[0,1]$ without loss of generality since linear rescaling of $Y$ does not fundamentally change the problem. Therefore, $\eta=\sigma^{2}/2\in[0,1/2]$. In this case, the rate in Theorem 2 contracts faster than $(1/2)^{K+1}$ as $K$ increases.
Remark 3.
[…]
Proof of Theorem 2.
Recall Lemma 4 and the definition (2.4), we have
[…]
From Lemma 5,
@@
Here the constant $\tilde{C}$ is a universal constant that does not depend on $\eta$ nor $y$; it only depends on $\|F^{-1}\|_{\mathrm{Lip}}$ and $\|F^{(k)}\|_{\infty}$ for $k\leq 2K+2$. ∎
@@ 3 Higher-Order Score Denoising: Identification and Estimation
The previous section establishes the noise asymptotic expansion of the optimal transport map pushing $Q$ to $P$. However, the expansion in Theorem 1 depends on the higher-order score functions of $P$, unknown in practice. This section presents an alternative approach that quantifies the denoising expansion in terms of the higher-order score functions of $Q$, which is identifiable and estimable from the noisy measurements $Y$.
@@
Definition 4 (Hölder Smooth Functions).
@@
$\mathcal{H}_{\Omega}^{\alpha}(L):=\left\{f:\Omega\to\mathbb{R}~\mid~\max_{0\leq k\leq\lfloor\alpha\rfloor}\sup_{y\in\Omega}|f^{(k)}(x)|+\sup_{x\neq y\in\Omega}\frac{|f^{(\lfloor\alpha\rfloor)}(x)-f^{(\lfloor\alpha\rfloor)}(y)|}{|x-y|^{\alpha-\lfloor\alpha\rfloor}}\leq L\right\}\;.$

### v3

@@ 1.3 Relations to the Literature
The closest works to ours are [GS24, JIS25, Lia25], which study denoisers motivated by OT theory [AGS05, Vil08]. Motivated by the fact that Bayes-optimal and empirical Bayes denoisers typically shrink the distribution too aggressively, these three papers propose different remedies. [GS24] and [JIS25] assume the existence of certain structural information of the prior distribution $P$, and solve for denoisers that minimize the mean squared error $\mathop{\mathbb{E}}\|T(Y)-X\|^{2}$ subject to such distributional constraints, say $T\sharp Q=P$; one can also view them as post-processing empirical Bayes denoisers to avoid over-shrinkage for better distributional matching.
In contrast, this paper and [Lia25] do not require any a priori knowledge of $P$, and construct optimal transport-inspired denoisers solely based on the noisy measurements $Y\sim Q$. This paper generalizes the results in [Lia25] to arbitrary higher-order denoisers $T_{K},K\geq 1$ in the univariate setting. This paper characterizes the fundamental role of higher-order score functions in constructing the hierarchy of OT denoisers $T_{1},T_{2},\ldots,T_{\infty}$, and reveals the combinatorial structure underlying the infinite expansion of the optimal transport map via Bell polynomial recursions.
1.4 Notation
@@
We use the following notations throughout the paper. We use lower-case letters such as $x,y,z$ to denote scalars in $\mathbb{R}$, and upper-case letters such as $X,Y,Z$ to denote random variables in $\mathbb{R}$. We reserve $f,g,h:\mathbb{R}\to\mathbb{R}$ and $F,G,T:\mathbb{R}\to\mathbb{R}$ to denote measurable functions or maps. $P,Q$ denote probability distributions on $\mathbb{R}$, with density functions $p,q:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ respectively with respect to the Lebesgue measure; $\phi:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ denotes the standard normal density function. $\sigma>0$ denotes the noise level, and we call $\eta=\sigma^{2}/2$ the noise parameter.
@@
For a smooth function $f:\mathbb{R}\to\mathbb{R}$ and an integer $k$, we use $f^{(k)}$ to denote the $k$-th order derivative of $f$. For a probability distribution $P$ on $\mathbb{R}$ with density $p$, we use $\mathcal{L}(X)$ to denote the distribution of the random variable $X\sim P$. We use $T\sharp Q$ to denote the push-forward distribution of $Q$ via a measurable map $T:\mathbb{R}\to\mathbb{R}$, i.e., for any measurable set $A\subseteq\mathbb{R}$, $(T\sharp Q)(A)=P_{Y\sim Q}(T(Y)\in A)$. For two probability distributions $P,Q$ on $\mathbb{R}$, we use $W_{r}(P,Q)$ to denote the $r$-Wasserstein distance between $P$ and $Q$ with $r\geq 1$. Given a function $f$ and a probability distribution $Q$, we denote the $L^{2}(Q)$-norm as $\|f\|_{L^{2}(Q)}=\left(\mathop{\mathbb{E}}_{Y\sim Q}f(Y)^{2}\right)^{1/2}$. We use $\|f\|_{\infty}=\sup_{x}|f(x)|$ to denote the $L^{\infty}$-norm of $f$, $\|f\|_{\mathrm{Lip}}$ to denote the uniform Lipschitz constant for $f$.
@@
Given i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$, we denote the empirical expectation $\widehat{\mathop{\mathbb{E}}}_{n}f=\tfrac{1}{n}\sum_{i=1}^{n}f(Y_{i})$. $\mathcal{F},\mathcal{H}$ denote function classes. We use $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|)$ to denote the covering number of a function class $\mathcal{F}$ at scale $\epsilon$ under the norm $\|\cdot\|$. We use the standard asymptotic notations $f(n)=O(g(n))$, or equivalently $f(n)\precsim g(n)$, if there exists a constant $C>0$ such that $|f(n)|\leq C|g(n)|$ for all sufficiently large $n$; we write $f(n)\asymp g(n)$ if $f(n)\precsim g(n)$ and $g(n)\precsim f(n)$.
@@ 2 Optimal Transport Expansion and Noise Asymptotics
We note that it is without loss of generality to assume $\sigma\in[0,1]$ and thus $\eta=\sigma^{2}/2\in[0,1/2]$: denote $\gamma^{2}:=\mathrm{Var}[Y]$, then we can rescale $Y$ to $\widetilde{Y}=Y/\gamma=\widetilde{X}+\widetilde{\sigma}Z$ where $\widetilde{X}=X/\gamma$ and $\widetilde{\sigma}=\sigma/\gamma\in[0,1]$. It follows that the noise level $\sigma$ can always be normalized to be at most $1$ by rescaling the data.
@@
Definition 2.1 (Model).
Let $X\in\mathbb{R}$ be a random variable with distribution $P$ and a valid density $p$ with respect to the Lebesgue measure. Let $Z\sim\mathcal{N}(0,1)$ be independent standard normal noise. Consider the noisy measurement $Y=X+\sigma Z$ with distribution $Q$ and density $q$. Let $\eta=\sigma^{2}/2$.
@@
$F(x)=\int_{-\infty}^{x}p(z)\,\mathrm{d}z\;,\qquad G(y)=\int_{-\infty}^{y}q(z)\,\mathrm{d}z\;.$
The notations $F^{(k)}$ and $G^{(k)}$ denote the $k$-th order derivatives of $F,G$ respectively. $F^{-1},G^{-1}:[0,1]\to\mathbb{R}$ denote the quantile functions (generalized inverse) of $F$ and $G$ respectively, defined as
@@ 2.1 F-expansion of the Optimal Transport Map
Definition 2.2 (Optimal Transport Map).
@@
Such a map exists and is unique almost everywhere when $P,Q$ are absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}$ [Vil03, Theorem 2.18], and minimizes the $r$-Wasserstein distance between $P$ and $Q$ for any $r\geq 1$,
$W_{r}(P,Q)=\inf_{T:T\sharp Q=P}~\left(\int_{\mathbb{R}}|T(y)-y|^{r}q(y)\,\mathrm{d}y\right)^{1/r}\;.$
@@
Definition 2.3 (Bell Polynomials).
@@
$B_{n,k}(x_{1},x_{2},\ldots,x_{n-k+1})=\sum\frac{n!}{j_{1}!j_{2}!\cdots j_{n-k+1}!}\prod_{\ell=1}^{n-k+1}\left(\frac{x_{\ell}}{\ell!}\right)^{j_{\ell}}\;,$
@@
Theorem 2.4 (F-expansion of the Optimal Transport Map).
Consider the additive Gaussian noise model in Definition 2.1. Assume that $F$ is infinitely differentiable. Define the following series expansion
@@
where $g_{1},g_{2},\ldots:\mathbb{R}\rightarrow\mathbb{R}$ are functions defined iteratively through the Bell polynomials $B_{n,k}$ in Definition 2.3:
@@
Then, for $y\in\mathbb{R}$ such that $T_{F}(y)$ is absolutely convergent, $T_{F}$ represents the optimal transport map pushing $Q$ onto $P$ as in Definition 2.2, namely,
@@
    Equation (2.4) reads $g_{2}=-\frac{F^{(2)}}{F^{(1)}}B_{2,2}(g_{1})+\frac{F^{(4)}}{F^{(1)}}$, and thus we have
@@
    Equation (2.4) reads $g_{3}=-\frac{F^{(2)}}{F^{(1)}}B_{3,2}(g_{1},g_{2})-\frac{F^{(3)}}{F^{(1)}}B_{3,3}(g_{1})+\frac{F^{(6)}}{F^{(1)}}$, and thus we have
@@
    $B_{4,1}=g_{4},\quad B_{4,2}=4g_{1}g_{3}+3g_{2}^{2},\quad B_{4,3}=6g_{1}^{2}g_{2},\quad B_{4,4}=g_{1}^{4}.$
    Equation (2.4) reads $g_{4}=-\tfrac{F^{(2)}}{F^{(1)}}B_{4,2}-\tfrac{F^{(3)}}{F^{(1)}}B_{4,3}-\tfrac{F^{(4)}}{F^{(1)}}B_{4,4}+\tfrac{F^{(8)}}{F^{(1)}}$, and thus we have
@@ 2.2 Higher-Order Noise Asymptotics
In this section, we derive the accuracy of higher-order denoisers based on the noise asymptotic expansion of the optimal transport map in Theorem 2.4. Define the $K$-th order denoiser $T_{K}:\mathrm{supp}(P)\to\mathbb{R}$ as the truncated series expansion of $T$ up to the $K$-th order term,
@@
Since $\{g_{i}\}_{i=1}^{K}$ depend on higher-order score functions $\tfrac{F^{(k)}}{F^{(1)}},k\leq 2K$, which are well-defined on the support of $P$, we concern ourselves only with the behavior of $T_{K}$ on the support of $P$.
@@
$W_{r}^{\ast}(T\sharp Q,P):=\left(\int_{\mathrm{supp}(P)}\left|T(y)-F^{-1}\circ G(y)\right|^{r}q(y)\,\mathrm{d}y\right)^{1/r}\;.$ (2.4)
In the case when $\mathrm{supp}(P)=\mathbb{R}$, this definition coincides with the standard $r$-Wasserstein distance $W_{r}(T\sharp Q,P)$; see Lemma A.4 for this equivalence.
The next theorem studies the $W_{r}^{\ast}(\cdot,\cdot)$ between the denoised distribution $T_{K}\sharp Q$ and the signal distribution $P$, as well as the uniform approximation error between the denoiser $T_{K}$ and the optimal transport map $F^{-1}\circ G$ restricted to the support of $P$.
Theorem 2.6 (Higher-Order Accuracy).
Let $K\geq 1$ and consider the additive Gaussian noise model in Definition 2.1 with $F$ that is $(2K+2)$-differentiable. Let $T_{K}$ be the $K$-th order denoiser defined above with $\{g_{i}\}_{i=1}^{K}$ defined in Theorem 2.4. Assume that $\mathrm{supp}(P)$ is compact and $\inf_{x\in\mathrm{supp}(P)}F^{(1)}(x)>0$. Recall (2.4), for any $r\geq 1$, we have
@@
Remark 2.7.
As noted earlier, we can assume that $\sigma\in[0,1]$ without loss of generality since linear rescaling of $Y$ does not fundamentally change the problem. Therefore, $\eta=\sigma^{2}/2\in[0,1/2]$. In this case, the rate in Theorem 2.6 contracts faster than $(1/2)^{K+1}$ as $K$ increases.
Remark 2.8.
We note that $T_{K}$ is monotonically increasing when $\eta$ is small. Indeed, since $T_{K}$ is a finite sum of smooth functions, it suffices to show that the first derivative $\frac{\mathrm{d}y}{\mathrm{d}T}_{K}(y)$ is positive when $\eta$ is small enough. Note that
$\frac{\mathrm{d}y}{\mathrm{d}T}_{K}(y)=1+\sum_{k=1}^{K}\frac{\eta^{k}}{k!}\frac{\mathrm{d}y}{\mathrm{d}g}_{k}(y)\;.$
Therefore, when $\eta$ is sufficiently small, $\frac{\mathrm{d}y}{\mathrm{d}T}_{K}(y)>0$ for all $y\in\mathrm{supp}(P)$ compactly supported.
Proof of Theorem 2.6.
By the definition (2.4), we have
$W_{r}^{\ast}\big(T_{K}\sharp Q,P)$ $=\left(\int_{\mathrm{supp}(P)}\left|T_{K}(y)-F^{-1}\circ G(y)\right|^{r}q(y)\,\mathrm{d}y\right)^{1/r}\;.$
From Lemma A.5,
@@
Here the constant $\tilde{C}$ is a universal constant that does not depend on $\eta$ or $y$; it only depends on $\|F^{-1}\|_{\mathrm{Lip}}$ and $\|F^{(k)}\|_{\infty}$ for $k\leq 2K+2$. ∎
@@ 3 Higher-Order Score Denoising: Identification and Estimation
The previous section establishes the noise-asymptotic expansion of the optimal transport map pushing $Q$ to $P$. However, the expansion in Theorem 2.4 depends on the higher-order score functions of $P$, unknown in practice. This section presents an alternative approach that quantifies the denoising expansion in terms of the higher-order score functions of $Q$, which are identifiable and estimable from the noisy measurements $Y$.
@@
Definition 3.1 (Hölder Smooth Functions).
@@
$\mathcal{H}_{\Omega}^{\alpha}(L):=\left\{f:\Omega\to\mathbb{R}~\mid~\max_{0\leq k\leq\lfloor\alpha\rfloor}\sup_{x\in\Omega}|f^{(k)}(x)|+\sup_{x\neq y\in\Omega}\frac{|f^{(\lfloor\alpha\rfloor)}(x)-f^{(\lfloor\alpha\rfloor)}(y)|}{|x-y|^{\alpha-\lfloor\alpha\rfloor}}\leq L\right\}\;.$